OpenStudy (anonymous):

Test for symmetry with respect to theta=pie/2 , the polar axis, and the pole. with r=16cos3theta

OpenStudy (anonymous):

\[r=16 \cos 3 \theta\] and \[\theta=\pi/2\] Then \[r=16 \cos (3 \pi/2) = 0\] Since \[\cos (-\theta) = \cos \theta \] \[r=16 \cos (-3 \pi/2) = 0\] Therefore, with respect to polar axis, the equation is symmetric about pole.